# Learning Exhaustive Correlation for Spectral Super-Resolution: Where Unified Spatial-Spectral Attention Meets Mutual Linear Dependence

### 20 Dec 2023 v1

@@ 4.3.1 Quantitative Results
On NTIRE 2022 [5], NTIRE 2020 [4], and ICVL [3] datasets, we compared ECT with various neural networks, as presented in Table 1. HSCNN+ is the champion in the NTIRE 2018 Clean Track and Real World Track. HR Net and AWAN are the champions in the NTIRE 2020 Clean track and Real World track, respectively. MST++ and DRCR are the first and third place in NTIRE 2022, respectively. Restormer is an advanced algorithm in image reconstruction. MST-L is a CASSI-based HSI reconstruction algorithm. Among all the methods, ECT achieves the lowest MRAE with the lowest computational costs and the smallest number of parameters, demonstrating the superior performance of our method, and highlighting the significance of modeling unified spatial-spectral correlation and mutual linear dependence.
@@ 4.3.2 Qualitative Results
We showcase the visual effects of the MRAE heatmaps in Figure 6 and Figure 7. We also present a comparison of spectral curves in small regions among the Ground Truth and various reconstruction algorithms in Figure 8. The visual results indicate that ECT exhibits the best reconstruction performance across different wavelengths, which further confirms the effectiveness of our approach.
@@ 4.4 Results on Real Data
| Methods | Outdoor Scene | Indoor Scene | ||
| MRAE | RMSE | MRAE | RMSE | |
| DRCR [25] | 0.2143 | 0.0070 | 0.2499 | 0.0104 |
| MST++ [9] | 0.2622 | 0.0084 | 0.2257 | 0.0091 |
| ECT (Ours) | 0.2120 | 0.0065 | 0.2095 | 0.0081 |
To further investigate the generalization ability of ECT, we conduct some experiments on real RGB data. We capture RGB images of a color chart both indoors under halogen lights and outdoors under sunlight, along with corresponding HSIs. We evaluated the algorithms using the average error of the 18 color patches on the color chart. The quantitative and qualitative results of ECT compared with two advanced spectral super-resolution methods MST++ [9] and DRCR [25] are shown in Table 2 and Figure 9. The experimental results on real data demonstrate the advanced performance and the strong generalization ability of our ECT.
4.5 Ablation Study
To fully explore the effects and the working mechanics of the Spectral-wise Discontinuous 3D (SD3D) splitting strategy and the Dynamic Low-Rank Mapping (DLRM) module, we introduce some ablation studies here. More ablation studies can be found in the supplementary material. All ablation studies are conducted on the NTIRE 2022 dataset with a one-stage U-shaped net.
| SD3D | DLRM | MRAE | RMSE | Params |
FLOPs |
| ✗ | ✗ | 0.1761 | 0.0266 | 0.55M |
7.84G |
| ✓ | ✗ | 0.1700 | 0.0255 | 0.58M |
8.24G |
| ✗ | ✓ | 0.1733 | 0.0261 | 0.56M |
8.27G |
| ✓ | ✓ | 0.1665 | 0.0246 | 0.60M |
8.69G |
4.5.1 Ablation Study of SD3D Splitting and DLRM
The Spectral-wise Discontinuous 3D (SD3D) splitting strategy and the Dynamic Low-Rank Mapping (DLRM) are our two significant contributions. The SD3D splitting strategy is used to model unified spatial-spectral correlation, while DLRM is employed to capture mutual linear dependence. The experimental results, as shown in Table 3, demonstrate that both the SD3D splitting strategy and DLRM can lead to improvements. When used together, they achieve even greater performance improvements.
@@ 4.5.1 Ablation Study of SD3D Splitting and DLRM
4.5.2 Ablation Study of SD3D Splitting Strategy
The key characteristic of the SD3D splitting strategy lies in its spatial-wise continuity and spectral-wise discontinuity splitting approach. Discontinuous splitting allows for a greater focus on non-local information, while continuous splitting helps preserve the local structure. We conducted an ablation study on the continuity and discontinuity in both spectral and spatial directions, as shown in Table 4. Experiments indicate that spectral super-resolution benefits from non-local features in the spectral direction, and adverse effects arise when disrupting spatial continuity.
4.5.3 Ablation Study of Low-Rank Factor
The critical parameter in the DLRM module is the number of feature () columns, denoted as . The rank of the low-rank dependence map in DLRM is not greater than . The experimental results for different values of in ESAB are shown in Table 5. When , it means , which does not constrain the dependence map to be low-rank. The experimental results highlight the importance of the low-rank characteristic of the dependence map.
| Fac. | MRAE | RMSE | Params | FLOPs |
@@ 5 Conclusion
In this paper, we analyze the limitations of existing spectral super-resolution Transformers in modeling unified spatial-spectral correlation and mutual linear dependence. We propose a novel Exhaustive Correlation Transformer (ECT) to model these correlations for spectral super-resolution. Specifically, we propose a Spectral-wise Discontinuous 3D (SD3D) splitting strategy to model unified spatial-spectral correlation and a Dynamic Low-Rank Mapping (DLRM) module to capture mutual linear dependence. Experimental results demonstrate that our approach achieves state-of-the-art performance on both simulated and real data. In future work, we will commit to applying our method to other tasks related to hyperspectral imaging.

### 18 Mar 2024 v2

@@ 4.3.1 Quantitative Results
On NTIRE 2022 [5], NTIRE 2020 [4], and ICVL [3] datasets, we compared ECT with various neural networks, as presented in Table 1. HSCNN+ is the champion in the NTIRE 2018 Clean Track and Real World Track. HR Net and AWAN are the champions in the NTIRE 2020 Clean track and Real World track, respectively. MST++ and DRCR are the first and third place in NTIRE 2022, respectively. Restormer is an advanced algorithm in image reconstruction that has a core design similar to MST++. HySAT is the SOTA spectral super-resolution method published recently. Among all the methods, ECT achieves the lowest MRAE with the lowest computational costs and the smallest number of parameters. We further test the inference latency of all models with the same input size on the same 3090 GPU. We select the average inference latency over 30 runs when the inference latency is stabilized. It is worth noting that compared to the existing SOTA HySAT, our ECT further enjoys a 34% reduction in inference time. The results demonstrate the superior performance of our method and highlight the significance of modeling unified spatial-spectral correlation and linear dependence.
| Method | Outdoor Scene | Indoor Scene | ||||
| MRAE | RMSE | SAM | MRAE | RMSE | SAM | |
| DRCR | 0.2143 | 0.0070 | 0.1896 | 0.2499 | 0.0104 | 0.2340 |
| MST++ | 0.2622 | 0.0084 | 0.2245 | 0.2257 | 0.0091 | 0.2055 |
| HySAT | 0.2135 | 0.0070 | 0.1868 | 0.2202 | 0.0088 | 0.1974 |
| ECT | 0.2012 | 0.0065 | 0.1730 | 0.2114 | 0.0082 | 0.1831 |
| MRAE | RMSE | Params | Latency | |
| 1 | 0.1648 | 0.0243 | 0.60 M | 41 ms |
| 2 | 0.1564 | 0.0236 | 1.19 M | 82 ms |
| 3 | 0.1542 | 0.0231 | 1.78 M | 124 ms |
@@ 4.3.2 Qualitative Results
We showcase the visual effects of the MRAE heatmaps in Figure 6, Figure 7, and Figure 8. We also present a comparison of spectral curves in small regions among the Ground Truth and various reconstruction algorithms in Figure 9. The visual results indicate that ECT exhibits the best reconstruction performance across different wavelengths, which further confirms the effectiveness of our approach.
@@ 4.4 Results on Real Data
To further investigate the generalization ability of ECT, we conduct some experiments on real RGB data. We choose flattened regions for validation to avoid the influence of misalignment. We capture RGB images of a color chart both indoors under halogen lights and outdoors under sunlight, along with corresponding HSIs. We evaluated the algorithms using the average error of the 18 color patches on the color chart. We further evaluate Spectral Angle Mapper (SAM) on real data. The quantitative and qualitative results of ECT compared with three advanced spectral super-resolution methods MST++ [9], DRCR [25] and HySAT [41] are shown in Table 4.3.1 and Figure 10. The experimental results on real data demonstrate the advanced performance and the strong generalization ability of ECT.
| SD3D | DLRM | MRAE | RMSE | Params | FLOPs |
| ✗ | ✗ | 0.1761 | 0.0266 | 0.55 M | 7.84 G |
| ✓ | ✗ | 0.1700 | 0.0255 | 0.58 M | 8.24 G |
| ✗ | ✓ | 0.1733 | 0.0261 | 0.56 M | 8.27 G |
| ✓ | ✓ | 0.1648 | 0.0243 | 0.60 M | 8.69 G |
| Splitting | MRAE | RMSE | Params | FLOPs |
| Spectral-wise | 0.1740 | 0.0257 | 0.59 M | 8.69 G |
| Spatial-wise | 0.1937 | 0.0285 | 0.94 M | 14.35 G |
| SD3D | 0.1648 | 0.0243 | 0.60 M | 8.69 G |
@@
| Factor | MRAE | RMSE | Params | FLOPs |
@@
4.5 Ablation Study
To fully explore the effects and the working mechanics of the whole architecture, Spectral-wise Discontinuous 3D (SD3D) splitting strategy, and the Dynamic Low-Rank Mapping (DLRM) module, we introduce some ablation studies here. More ablation studies can be found in the supplementary material. All ablation studies are conducted on the NTIRE 2022 dataset.
4.5.1 Ablation Study of the Network Stage
We conduct an ablation study on the number of stages () in the network, and the results are shown in Table 4.3.1. We primarily set balancing performance and inference latency. All ablation studies below are conducted using a 1-stage structure to efficiently validate the effectiveness of each module.
4.5.2 Ablation Study of SD3D Splitting and DLRM
The Spectral-wise Discontinuous 3D (SD3D) splitting strategy and the Dynamic Low-Rank Mapping (DLRM) are our two significant contributions. The SD3D splitting strategy is used to model unified spatial-spectral correlation, while DLRM is employed to capture linear dependence. We test the performance improvement of SD3D and DLRM compared to MST++ [9] without these two structures. The experimental results, as shown in Table 4.4, demonstrate that both the SD3D splitting strategy and DLRM can lead to improvements. When used together, they achieve even greater performance improvements. The results indicate the effectiveness of our two key designs.
4.5.3 Ablation Study of SD3D Splitting Strategy
First, We compare our SD3D splitting strategy with traditional spatial-wise and spectral-wise splitting strategies. The results shown in Table 4.4 demonstrate the effectiveness of our method. Moreover, the key characteristic of the SD3D splitting strategy lies in its spatial-wise continuity and spectral-wise discontinuity splitting approach. Discontinuous splitting allows for a greater focus on non-local information, while continuous splitting helps preserve the local structure. We conducted an ablation study on the continuity and discontinuity in both spectral and spatial directions, as shown in Table 4.4. Experiments indicate that spectral super-resolution benefits from non-local features in the spectral direction, and adverse effects arise when disrupting spatial continuity.
4.5.4 Ablation Study of Low-Rank Factor
The critical parameter in the DLRM module is the number of feature () columns, denoted as . The rank of the low-rank dependence map in DLRM is not greater than . The experimental results for different values of in ESAB are shown in Table 4.4. When , it means , which does not constrain the dependence map to be low-rank. The experimental results highlight the importance of the low-rank characteristic of the dependence map.
| Method | NTIRE 2022 | NTIRE 2020 | ICVL | Params (M) | FLOPs (G) | ||||||
| MRAE | RMSE | PSNR | MRAE | RMSE | PSNR | MRAE | RMSE | PSNR | |||
| PADUT | 0.1850 | 0.0271 | 33.04 | 0.0624 | 0.0158 | 37.07 | 0.0798 | 0.0193 | 36.77 | 1.71 | 20.29 |
| ECT | 0.1564 | 0.0236 | 34.81 | 0.0588 | 0.0142 | 37.71 | 0.0635 | 0.0142 | 38.50 | 1.19 | 16.75 |
5 Discussion
Recently, spectral reconstruction algorithms based on CASSI have received wide interest. We find that performance improvements brought by recent advanced CASSI-based spectral reconstruction algorithms [7, 20, 27] are typically reflected in sharper spatial reconstructions. However, spatial details reconstruction is not a necessary challenge in the spectral super-resolution task. Therefore, recent algorithms used to improve CASSI-based spectral reconstruction have difficulty enhancing the performance of spectral super-resolution. We retrain the SOTA algorithm for CASSI-based spectral reconstruction PADUT [27] for the spectral super-resolution task. The experimental results shown in Table 8 demonstrate that our ECT achieves significantly better performance.
@@ 6 Conclusion
In this paper, we analyze the limitations of existing spectral super-resolution Transformers in modeling unified spatial-spectral correlation and linear dependence. We propose a novel Exhaustive Correlation Transformer (ECT) to model these correlations for spectral super-resolution. Specifically, we propose a Spectral-wise Discontinuous 3D (SD3D) splitting strategy to model unified spatial-spectral correlation and a Dynamic Low-Rank Mapping (DLRM) module to capture linear dependence. Experimental results demonstrate that our approach achieves state-of-the-art performance on both simulated and real data with a 34% reduction in inference latency. In future work, we will commit to applying our method to other tasks related to hyperspectral imaging.